OpenStudy (anonymous):

Ralph is 3 times as old as Sara. In 6 years, Ralph will be only twice as old as Sara will be then. Find Ralph's age now.

OpenStudy (mww):

Best to set up an equation to describe the information. Let R be Ralph's CURRENT age and S be Sara's current age. If he is 3 times as old as Sara, then R = 3S (i) In 6 years, Ralph's age will be R+6. In 6 years Sara's age will be S+6 So R+6 = 2(S+6) as Ralph's age then will be twice Sara's age then. That's why we need to add 6 to both their ages as 6 yrs have passed. So we have two equations to solve: \[R=3S \] and \[R+6 =2(S+6) = 2S +12\] Now you need to simultaneously solve them either with substitution or elimination method to find R.